OpenStudy (taylovesmatthew):

. One hundred nautical miles equals about 185 kilometers. To the nearest kilometer, how far in kilometers is 290 nautical miles? A. 537 km B. 157 km C. 185 km D. 53,650 km

OpenStudy (juicstice):

set up a proportion

OpenStudy (anonymous):

A.

OpenStudy (juicstice):

it's B

OpenStudy (taylovesmatthew):

i asked my friend she said c i don't know which it is..

OpenStudy (anonymous):



OpenStudy (anonymous):

Its A

OpenStudy (juicstice):

100/185=x/290 100x290=29000 29000÷185=156.75

OpenStudy (taylovesmatthew):

i understand now thank you

OpenStudy (anonymous):

290 times 185 = 53650 53650 divided by 100 = 536.5